x^y^x^y^x^y..... = y^x^y^x^y^x...... (^ means to the power) . then x,y can take which values?(i mean what are possible values of x&y)

Grade:12

1 Answers

Vikas TU
7 years ago
From the series, x^y^x^y^x^y..... => can be written =>as,
(x^y)^m = m
From the series,  y^x^y^x^y^x...... => can be written as:
(y^x)^n = n
m = n given therefore,
(x^y)^m = (y^x)^m
o
x^y = y^x
take log both sides,
ylogx = xlogy
y/x = logy/logx
Hence for all x= y values it would be satiffies accept x and y should be non – zero.